OpenStudy (anonymous):

10. Find all the zeroes of the equation. –3x4+ 27x2 + 1200 = 0 (2 points)

OpenStudy (anonymous):

can somone help plzz

OpenStudy (anonymous):

i dont get this i need help now pleaz

OpenStudy (anonymous):

All of the terms can be divided by -3, and then treated as a quadratic.

OpenStudy (anonymous):

can u show me step by step how im kinda confused pleaz

OpenStudy (anonymous):

What do you get after dividing all the terms by -3?

OpenStudy (anonymous):

-1x^4-8x^2+400=0

OpenStudy (anonymous):

i think i did it wrong

OpenStudy (anonymous):

Yeah, that's not quite it. \(\large –3x^4+ 27x^2 + 1200 = 0 ÷ -3 \rightarrow x^4-9x^2 -400 = 0 \)

OpenStudy (anonymous):

Now make a substitution of y=x^2

OpenStudy (anonymous):

i dont get what ur talking about could u be more spisi

OpenStudy (anonymous):

pleaz

OpenStudy (anonymous):

No, it is not possible for me to be more specific. I told you exactly what you need to do. Once you have it in quadratic form, you may factor it.

OpenStudy (anonymous):

i seriously dont get these pleaz help

OpenStudy (anonymous):

If you had the equation y^2−9y−400=0, would you be able to solve it?

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

wait yes i would

OpenStudy (anonymous):

Can you help me I dont understand how to do it at all :/

OpenStudy (anonymous):

i still did not get it

OpenStudy (anonymous):

Dang, I'm 14 lessons behind and I'm failing :/ I'm trying to get caught up and passing so I wont fail the semester

OpenStudy (anonymous):

@KenaRolin same here :(

OpenStudy (anonymous):

@KenaRolin @Natalovesya Samee :c But, I'm 27 days behind

OpenStudy (anonymous):

in Algebra 2, right? Cuz I'm 19 lessons behind in math :(

OpenStudy (anonymous):

Yeess, maybe you could help me catch up and then I could help you with the rest? Or any other class we might have together? I finished with most of mine except math

OpenStudy (anonymous):

Yeah same. ok! what lesson are you on? And what other classes do you have?

OpenStudy (anonymous):

Unit 5 Lesson 9. I have 3D Art Modeling, Biology, English 11, Psychology, and US History

OpenStudy (anonymous):

@Natalovesya Nevermind, I'm on the Unit 5 Test now. If you have done it, it would be GREAT if you could please, help me out. Thanks xx

OpenStudy (anonymous):

We have a loot of the same classes :) I also have Biology, Psychology, and US History

OpenStudy (anonymous):

Unit 5 Test? Isn't that after the Unit Portfolio? Cuz, I'm stuck on that :(

OpenStudy (anonymous):

@justagirll

OpenStudy (anonymous):

No, it was after a a quiz.

OpenStudy (anonymous):

Are you in Honors? @justagirll